candygirl220:

A health club charges a one-time initiation fee and a monthly fee. John paid 100 dollars for 2 months of membership. However, Peter paid 200 for 6 months of membership. How much will Sylvia pay for 1 year of membership?

umm:

@dude- This is a trick question e.e

Ultrilliam:

@563blackghost @Shadow @dude

sillybilly123:

A method: \(i = \text{ one-time initiation fee}\) \(m = \text{ monthly fee}\) \(J = 100 = i + 2m\) \(P = 200 = i + 6m\) \(P - J = 100 = 4m \implies m = \dots\) \(J = 100 \implies 100 = i + 2(\dots) \implies i = \dots\) Then for yer answer: \(S = i + 12 m = 50 + 12(\dots) = \dots\) And maybe do a result check: \(S - P = 6m = \dots \) [350]

sillybilly123:

make sense?

candygirl220:

yes

sillybilly123:

perfect :)